OpenStudy (anonymous):

MEDAL GIVEN TO CORRECT ANSWER. You kick a ball with a speed of 14 m/s at an angle of 51°. How far away does the ball land? A. 1.8 m B. 3.96 m C. 21.3 m D. 19.6 m

OpenStudy (anonymous):

Find the vertical velocity from vsuby = 14 sin (51) = 10.9 m/s Find how long until the velocity has become zero from vsuby = 10.9-(9.8) t = 0 due to the negative acceleration of gravity. t = 10.9/9.8=1.1 sec. You could calculate the height it reached d then calculate how long it takes to come back to ground, but the situation is symmetric, so it takes another 1.1 s to fall. total time is 2.2 s. Meanwhile, during those 2.2 s, the ball has had a horizontal component of velocity of vsubx = 14 cos(51= 8.8 m/s and goes (8.8 m/s)(2.2s) = 19.4 m . Looks like D, with a little difference from rounding.